OpenStudy (anonymous):

What parts do I need to integrate sin^(3x)*cos^2(x)dx? (Do I need to use algebraic substitution? If so what is the u value?)

zepdrix (zepdrix):

Did you mean to type \(\Large\rm \sin^3x\cos^2x\) ?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

If you use your pythagorean identity, you can clean this up a little bit. \[\Large\rm \sin x \sin^2x \cos^2x=\sin x(1-\cos^2x) \cos^2x\] u= cos x \[\Large\rm =-\int\limits (1-\cos^2x)\cos^2x (-\sin x~dx)\]

zepdrix (zepdrix):

\[\Large\rm =-\int\limits\limits (1-u^2)u^2 (du)\]Understand what I did there? Or was the thing with the negatives confusing?

OpenStudy (anonymous):

Nope, the negatives were fine! I was trying the wrong u-value. Thank you so much!!

zepdrix (zepdrix):

cool c:

OpenStudy (anonymous):

If you were wondering about a method using integration by parts, you could split up the powers like zep did: \[\int\sin^3x\cos^2x~dx=\int\sin^3 x\cos x\cos x~dx\] Then set the following: \[\begin{matrix} u=\cos x&&&dv=\sin^3x\cos x~dx\\ du=-\sin x~dx&&&dv=t^3~dt&&\text{(where }t=\sin x)\\ &&&v=\frac{1}{4}\sin ^4 x \end{matrix}\] which gives \[\int\sin^3x\cos^2x~dx=\frac{1}{4}\sin^4 x\cos x+\frac{1}{4}\int\sin^5x~dx\] The new integral can be handled with another power split and some identity usage: \[\int\sin x\sin^4x~dx=\int\sin x(1-\cos^2x)^2~dx=\int\sin x(1-2\cos^2x+\cos^4x)~dx\] and so on. This way involves a lot of power reduction, many more steps than zep's clearly superior method. Always check for substitutions before considering IBP.